The perimeter of an isosceles triangle is 40, and the base is 10. find the length of one of the equial sides.
Answer by mszlmb(115)

"The perimeter of an isosceles triangle is 40, and the base is 10. find the length of one of the equial sides."













Perimeter=Base+Side+Side

Thence;
40=B+S+S

We know S and S [two sides] are the same, because the triangle is an isosceles.
The equation thence yields:
40=B+2S

B is 10:
40=10+2S   minus 10 to both sides
30=2S      divide both sides by 2
15=S

each side is 15. 

let's check: 
2(15)+10=40
30+10=40
40=40 :-D it works.